# ScVLM: a Vision-Language Model for Driving Safety Critical Event Understanding

### v1

@@ 4.2 Application
Model Implementation For both classification tasks, the dataset was randomly divided into training, testing, and validation subsets in a 7:2:1 ratio. For the event type classification task, this corresponds to 11,981 events for training, 3,424 for testing, and 1,711 for validation. For the conflict type classification task, there are 6,030 SCEs for training, 1,724 for testing, and 861 for validation. The validation sets were used for hyper-parameter tuning, while evaluation performance was assessed on the independent testing sets. The software environment was based on Python 3.8 running on Rocky Linux 9.3. The model was trained on a high-performance GPU workstation equipped with dual Intel Xeon Gold 6338 CPUs @ 2.00 GHz, 256 GB RAM, and two Nvidia Tesla A100 80 GB GPUs.
@@
This study compares the performance of supervised learning and contrastive learning approaches on two classification tasks using five benchmarks. The selected contrastive learning algorithms are X-CLIP [23] and ActionCLIP [24], while the supervised learning algorithms include SlowFast [19], Video Swin Transformer [28], and TimeSformer [29]. These methods demonstrate competitive performance on the Kinetics-400 dataset [30], providing a suitable basis for evaluating the efficacy of the two learning approaches in two classification tasks.
@@
X-CLIP adopts the ViT-B/16 CLIP architecture as parts of the cross-frame communication transformer and a one-layer multi-frame integration transformer [23]. ActionCLIP utilizes the ViT-B/16 CLIP architecture with six Transformer adapter layers [24]. The SlowFast model employs a ResNet3D backbone structure [19]. The Video Swin Transformer is based on the Swin-Base model architecture [28]. The TimeSformer employs a TimeSformer-Base model with divided space-time attention [29]. For both TimeSformer and Video Swin Transformer, the weights are initialized from models pretrained on ImageNet-22k [31]. Each algorithm is trained using a batch size that maximizes the GPU memory capacity of two Tesla A100 GPUs. The validation set is used for evaluation after each epoch, and the epoch with the highest validation accuracy is selected for testing on an independent testing set.
For narrative generation, the VideoLLaMA2 model is employed, utilizing the CLIP ViT-Large-Patch14-336 as its video encoder and the Mistral-7B-Instruct-v0.2 as its language decoder [5]. The process begins by using the prompt “Describe this video” to extract static environment information. This information is then combined with classification results from the most effective models: SlowFast for event type classification and X-CLIP for conflict type classification. If the event type is “Normal Driving,” the narrative is generated with the prompt “Describe this event: 1: {Environment}. 2: Normal Driving.” For SCEs, the narrative is generated using the prompt “Describe this event: 1: {Environment}. 2: {Event Type}. 3: {Conflict Type}.”
4.3 Classification Task Performance
X-CLIP Contrastive 0.636 0.847 0.278 0.770 0.232 0.259 0.222
Action CLIP Contrastive 0.606 0.846 0.239 0.777 0.216 0.244 0.220
SlowFast Supervised 0.485 0.806 0.148 0.659 0.130 0.156 0.105
Swin Transformer Supervised 0.545 0.829 0.197 0.752 0.163 0.158 0.153
TimeSformer Supervised 0.571 0.840 0.205 0.786 0.166 0.177 0.154
Table 4: Comparison of different models in conflict type classification (Training Proportion = 0.05)
@@ 4.3 Classification Task Performance
Table 5: Comparison of alternative models in conflict type classification
Six evaluation metrics are employed to comprehensively assess the performance of the selected models. Accuracy measures the overall classification performance by using the ArgMax function to select the class with the highest predicted probability as the final output. Mean average precision (mAP) and the average area under the ROC curve (AUC) evaluate classification confidence across all thresholds. Balanced accuracy, macro precision, and macro F1 scores assess performance in imbalanced sample scenarios, emphasizing the rare-event nature of SCEs [32].
Benchmark Comparison Table 1 presents a comparative analysis of five benchmark models for four-way event type classification. All the supervised learning-based models outperformed the contrastive learning-based models in terms of accuracy, mAP, AUC, macro precision, and macro F1 score with SlowFast achieving the best performance across all evaluation metrics. This suggests that on the SHRP 2 NDS front-view video dataset, selected supervised learning approaches are more effective for event type classification task than the selected contrastive learning approaches.
Table 2 provides a comparative analysis of five benchmark models for 16-way conflict type classification. In this task, the contrastive learning-based models outperformed the supervised learning-based models in most evaluation metrics. Between the two contrastive learning-based models, X-CLIP demonstrates the best overall performance. This suggests that on the SHRP 2 NDS front-view video dataset, selected supervised learning approaches are less effective for conflict type classification than selected contrastive learning approaches.
Figure 6: Count of each conflict type with 5% training set.
Few-shot Classification Performance This experiment uses a randomly selected 5% of the training dataset to evaluate the performance of the two learning approaches in classifying conflict types under extremely limited data conditions. As shown in Figure 6, 10 out of the 16 conflict types have fewer than 10 occurrences within the 5% training dataset. By training the model on a smaller subset of data, we can assess its ability to make accurate predictions with minimal information. Table 4 provides the performance of five benchmark models for 16-way conflict type classification using 5% of the training dataset. In this limited data setting, the contrastive learning-based models significantly outperformed their supervised learning-based models across most evaluation metrics. The improvements were particularly notable in Balanced Accuracy, macro precision, and macro F1 score, indicating that the contrastive learning approach performs better on minority classes when data is scarce. This suggests that in the setting of data deficiency, selected supervised learning approaches are much less effective for conflict type classification than selected contrastive learning approaches.
Alternative model comparison To evaluate that the effective component of contrastive learning approaches in 16-way conflict type classification is the contrastive learning approach or the CLIP encoder, two alternative models are implemented. The raw video frames are processed through a CLIP image encoder, and two methods are employed to handle temporal dependencies across frames: mean pooling and long short-term memory (LSTM), similar to the methods used in ActionCLIP [24]. The video representation obtained from mean pooling or LSTM is then fed into a classifier, specifically a multi-layer perceptron, as shown in Figure 7.
@@
Table 5 compares X-CLIP and SlowFast, the leading models for conflict type classification task using contrastive learning and supervised learning respectively. The performance of the CLIP image encoder in the supervised learning approach is notably lower, suggesting that the CLIP image encoder may not be well-suited for supervised learning approach in conflict type classification. The superior performance of the CLIP-based contrastive learning approach is largely attributed to the model’s architecture, with the text encoder playing a crucial role, especially for labels with rich text.
4.4 Prompt Tuning for VLM
Figure 8: Prompt strategies
Hallucinations in VLM refer to the scenario that response towards the prompt are fabricated, inaccurate, or misleading, often arising from the model generating responses without sufficient factual support or context [33]. To mitigate these hallucinations, approaches to design and refine prompts for VLM are proposed, involving two key steps in prompt tuning: first, instructing the VLM to describe the static environment, and second, directing the VLM to generate a narrative description of SCEs [34].
@@ 4.4 Prompt Tuning for VLM
Chain-of-Thought prompt and the repetition of important information are two techniques utilized in prompt tuning [34, 35, 36]. In this study, a Chain-of Thought prompt combined with a repeat-answer strategy is employed. Initially, the VLM is prompted to describe the environment. Subsequently, the event type, conflict type, and the environment description are combined to prompt the VLM for the final narrative of the SCE, as shown in Figure 8 (c). To evaluate this approach, 20 SCEs are randomly selected from the testing set and assessed using three strategies: the proposed strategy and two alternatives. The first alternative, direct prompt (Figure 8 (a)), involves directly asking the VLM to describe the SCE based on the event type and conflict type. The second alternative, Chain-of-Thought prompt (Figure 8 (b)), involves first asking the model to describe the environment, followed by combining the event type and conflict type to generate the SCE narrative.
The narratives generated by three strategy were compared across 20 selected SCEs. The Chain-of-Thought combined with repeat-answer strategies yielded the most accurate descriptions with the fewest hallucinations. Three resulting narratives for a lead vehicle crash example on a highway are provided, with the key frames shown in Figure 9. The bold text highlights hallucinations.
@@
Figure 9: Key frames of a lead vehicle crash example
  • Direct Prompt: The video shows a car driving on a road, and suddenly, it crashes into the lead vehicle. The collision causes the car to stop abruptly, and the driver is seen trying to regain control of the vehicle. The car then moves forward, but it is hit again by the lead vehicle, causing it to stop once again. The driver is seen trying to move the car, but it is stuck in place. The video captures the entire sequence of the crash and the driver’s attempts to recover from the situation.
  • Chain-of-Thought Prompt: The video depicts a crash and conflict with the lead vehicle. The scene unfolds on a highway where a car is seen driving behind another vehicle. Suddenly, the car collides with the lead vehicle, resulting in a crash. The impact causes the lead vehicle to veer off course, and the car that caused the crash continues to drive ahead. The conflict between the two vehicles is evident as the car that caused the crash tries to overtake the lead vehicle, which is now struggling to maintain control on the road. The video captures the intense moment of the crash and the subsequent struggle for control between the two vehicles.
  • Chain-of-Thought and repeat-answer Prompt: The video depicts a crash scenario on a busy highway. Several cars are visible, moving along the wet road, which suggests recent rain or ongoing drizzle. The environment is urban, evidenced by the presence of tall buildings in the background. The sky is overcast with clouds, adding to the dreary atmosphere. White lines are clearly painted on the road, indicating lane divisions. The crash involves a conflict with the lead vehicle, implying that one vehicle has collided with another directly in front of it. The scene is typical of a congested highway setting where traffic density and road conditions may contribute to such incidents.
Both the direct prompt and Chain-of-Thought prompt frequently reference elements not present in the video. In contrast, the Chain-of-Thought and repeat-answer prompts perform best, with the only hallucination being the mention of a wet road. This issue may arise from sunlight reflections on the road that resemble wetness. Overall, the proposed Chain-of-Thought and repeat-answer prompt strategy provides the most accurate description of the SCE.
@@ 5 Conclusion
Based on the SHRP 2 NDS video dataset, the results demonstrate that the proposed ScVLM generates more precise and contextually appropriate event descriptions compared to a standard VLM. This work not only contributes to the accuracy of SCE detection but also offers a robust framework for future research in automatic generation of SCE descriptions.

### v2

@@ 4.2 Classification Task Implementation and Performance
Model Implementation The dataset was randomly split into training, testing, and validation subsets in a 7:2:1 ratio. Few-shot evaluation utilized 5% of the conflict type classification training set, with 10 categories containing fewer than 10 samples each. Validation sets were used for hyperparameter tuning, while independent testing sets assessed performance. The environment consisted of Python 3.8 on Rocky Linux 9.3, with model training performed on a workstation equipped with dual Intel Xeon Gold 6338 CPUs, 256 GB RAM, and two Nvidia Tesla A100 80 GB GPUs.
@@
This study evaluated five supervised and contrastive learning approaches to select a suitable method for ScVLM, including X-CLIP [24] and ActionCLIP [25] for contrastive learning, and SlowFast[20], Video Swin Transformer [21], and TimeSformer [29] for supervised learning. These five models have similar performance on the Kinetics-400 dataset [30].
@@
We used the following setup for each model. X-CLIP uses the ViT-B/16 CLIP architecture with a cross-frame communication transformer and a one-layer multi-frame integration transformer. ActionCLIP incorporates six Transformer adapter layers into the ViT-B/16 CLIP architecture. SlowFast employs a ResNet3D backbone, Video Swin Transformer utilizes the Swin-Base architecture, and TimeSformer adopts a TimeSformer-Base model with divided space-time attention. Both TimeSformer and Video Swin Transformer were initialized with ImageNet-22k pre-trained weights. All models were trained with batch sizes optimized for two Tesla A100 GPUs, with the best validation accuracy epoch selected for testing on an independent set.
@@
Table 4: Comparison of alternative models in conflict type classification.
Benchmark Comparison Six metrics were used to evaluate model performance: Accuracy, mean average precision (mAP), area under the ROC curve (AUC), balanced accuracy, macro precision, and macro F1. The latter three focus on imbalanced scenarios, suitable for the rare-event nature of SCEs [31].
Table 2 presents the results of four-way event type classification. In general, the supervised learning-based models outperformed the contrastive learning-based models, with SlowFast achieving the best performance across all evaluation metrics. This suggests that on the SHRP 2 NDS dataset, selected supervised learning approaches are more effective for event-type classification task than the selected contrastive learning approaches.
Table 3 presents a comprehensive comparison for 16-way conflict type classification. The results include both the full dataset and a limited 5% training subset for few-shot learning. In the full dataset analysis, contrastive learning-based models outperformed supervised learning-based models across most evaluation metrics, with X-CLIP demonstrating the best overall performance on the SHRP 2 NDS front-view video dataset.
For few-shot learning, contrastive learning-based models significantly outperformed their supervised learning counterparts across most metrics, with notable improvements in balanced accuracy, macro precision, and macro F1 score. These results indicate that contrastive learning approaches are more effective for conflict type classification than supervised learning approaches on the SHRP 2 NDS dataset, particularly when dealing with minority classes and limited data availability.
Component Performance Evaluation To evaluate whether the most effective component of contrastive learning approaches in 16-way conflict type classification was the contrastive learning approach or the CLIP encoder, two alternative models were implemented. The raw video frames were processed through a CLIP image encoder, and two methods were employed to handle temporal dependencies across frames: mean pooling and long short-term memory (LSTM) [25]. The resulting video representations from each method were input into a multi-layer perceptron classifier. The overall process is illustrated in Figure 6.
@@
Table 4 compares X-CLIP and SlowFast, the leading models for conflict type classification task using contrastive learning and supervised learning respectively. The performance of the CLIP image encoder in the supervised learning approach was notably lower, suggesting that the CLIP image encoder may not be well-suited for supervised learning approaches in conflict type classification. This evaluation confirms that the superior performance of the CLIP-based contrastive learning approach can be attributed to the model’s architecture, with the text encoder playing a crucial role, especially for labels with rich text.
4.3 Narrative Generation Implementation and Performance
Model Implementation The narrative generation process consists of two steps: environment information extraction and narrative generation. The VideoLLaMA2 model is employed for understanding environment information using the prompt “Describe this driving event from dashcam view." We used CLIP ViT-Large-Patch14-336 as the video encoder and Mistral-7B-Instruct-v0.2 as the language decoder [6].
Narrative generation combines generated environment information with classification results, based on the most effective models: SlowFast for event-type classification and X-CLIP for conflict-type classification. Narrative generation used LLaMA 3.1 8B [32] with the system prompt “This is related to a driving event. Describe objectively." If the event type was “Normal Driving," the narrative was generated with the user prompt “Describe this event: 1: Environment. 2: Normal Driving." For SCEs, the narrative was generated using the user prompt “Describe this event: 1: Environment. 2: Event Type. 3: Conflict Type." To make a fair comparison with other VLMs, the language models were not fine-tuned.
@@ 4.3 Narrative Generation Implementation and Performance
Alternative Prompts for Language Models Hallucinations in VLMs occur when responses lack factual support or context [33]. To mitigate this, this study employs a chain-of-thought prompt [34] combined with a repeat-answer [35] strategy. The VLM first describes the environment, then generates the SCE narrative using event type, conflict type, and environment description. To evaluate this approach, 20 randomly selected SCEs from the testing set were assessed using three strategies: 1) direct prompt 2) chain-of-thought prompt 3) chain-of-thought with repeat-answer (proposed strategy). These are illustrated in Figure 7.
Figure 7: Prompt strategies.
As an example, Figure 8 presents key frames and generated narratives for a lead vehicle crash on a highway, with hallucinations highlighted in red. Among the different strategies, the proposed approach produced the most accurate descriptions, minimizing hallucinations.
@@
Figure 8: Key frames and generated narratives for a lead vehicle crash example (red highlights are the hallucinations).
Benchmark Comparison for Narrative Generation To evaluate the performance of ScVLM, 100 events were randomly selected, and ground truth narratives were manually annotated by trained volunteers. The annotations include traffic density, light conditions, surface conditions, and locality. For SCEs, additional annotations specify event type, conflict type, and incident type. Two examples are provided, with the key frames shown in Figure 9.
Figure 9: Ground truth narratives for normal driving and SCE examples.
Among the 100 selected events, 56 were classified as SCEs. The analysis was conducted on two evaluation sets: the full set of selected events and the subset of SCEs. Ten state-of-the-art understanding VLMs were evaluated against ScVLM using metrics ROUGE-L [36], METEOR [37], and BERTScore [38]. The benchmark models used their default setup. For fair comparison, the benchmarks employed a chain-of-thought prompting approach with two sequential prompts: “Describe this driving event from dashcam view." and “If there is a safety critical event, describe it." The responses were then combined to form the final narrative. To comprehensively evaluate the generative narratives relative to the ground truth, F1 scores from ROUGE-L and BERTScore were used, providing a balanced measure of both precision and recall.
As shown in Table 5, ScVLM outperformed all other models in the full evaluation set, achieving the highest ROUGE-L and BERTScore. This demonstrates that ScVLM excels in narrative generation tasks compared to existing benchmarks. In the SCE subset, ScVLM shows a more pronounced advantage, surpassing all models across every evaluation metric. Specifically, ScVLM outperformed the second-best models by 13.7% in ROUGE-L, 7.5% in METEOR, and 5.0% in BERTScore. This substantial improvement highlights ScVLM’s superior performance in SCE narrative generation. ScVLM is the only model to show consistent improvement across all evaluation metrics in the SCE setting. This indicates that ScVLM is particularly robust to the challenges posed by the SCE scenario, outperforming all other models in terms of narrative generation quality.
Model& Eval set Size ROUGE-L METEOR BERT
ScVLM (all) 15B 0.186 0.197 0.572
VideoLLaMA 2 [6] (all) 7B 0.165 0.119 0.543
Qwen2-VL [39] (all) 72B 0.173 0.173 0.559
MiniCPM-V [40] (all) 8B 0.169 0.179 0.563
LLaVA-Next-Video [41] (all) 34B 0.138 0.204 0.542
LLaVA-OneVision [42] (all) 72B 0.159 0.163 0.525
Chat-UniVi [43] (all) 13B 0.162 0.199 0.559
LLaMA-Vid [44] (all) 13B 0.178 0.142 0.536
PLLaVA [45] (all) 13B 0.130 0.189 0.530
Video-ChatGPT [46] (all) 7B 0.132 0.145 0.521
Video-LLaVA [47] (all) 7B 0.161 0.167 0.541
ScVLM (SCE) 15B 0.199 $\uparrow$ 0.215 $\uparrow$ 0.589 $\uparrow$
VideoLLaMA 2 [6] (SCE) 7B 0.163 $\downarrow$ 0.120 $\uparrow$ 0.548 $\uparrow$
Qwen2-VL [39] (SCE) 72B 0.169 $\downarrow$ 0.169 $\downarrow$ 0.555 $\downarrow$
MiniCPM-V [40] (SCE) 8B 0.168 $\downarrow$ 0.179 -- 0.558 $\downarrow$
LLaVA-Next-Video [41] (SCE) 34B 0.135 $\downarrow$ 0.200 $\downarrow$ 0.533 $\downarrow$
LLaVA-OneVision [42] (SCE) 72B 0.157 $\downarrow$ 0.166 $\uparrow$ 0.525 --
Chat-UniVi [43] (SCE) 13B 0.157 $\downarrow$ 0.194 $\downarrow$ 0.561 $\uparrow$
LLaMA-Vid [44] (SCE) 13B 0.175 $\downarrow$ 0.145 $\uparrow$ 0.545 $\uparrow$
PLLaVA [45] (SCE) 13B 0.122 $\downarrow$ 0.177 $\downarrow$ 0.522 $\downarrow$
Video-ChatGPT [46] (SCE) 7B 0.121 $\downarrow$ 0.134 $\downarrow$ 0.520 $\downarrow$
Video-LLaVA [47] (SCE) 7B 0.154 $\downarrow$ 0.161 $\downarrow$ 0.538 $\downarrow$
Table 5: Comparison of different models in narrative generation.
@@ 5 Conclusion
Based on the SHRP 2 NDS video dataset, the results demonstrate that the proposed ScVLM generates more precise and contextually appropriate event descriptions compared to a standard VLM. This work not only contributes to the accuracy of SCE detection, but also offers a robust framework for future research in automatic generation of SCE descriptions.